Trey finished his chemistry assignment in 3/4 hours. Then he completed his history assignment in 5/8 hours. How much more time did Trey spend on his chemistry assignment?


1
Expert's answer
2020-08-25T13:06:09-0400

We calculate the difference between times spent on chemistry and history:

(3/4 - 5/8) hours = (3*2 -5)/8 hours= 1/8 hours = 60/8 minutes.